OpenStudy (anonymous):

How much money will there be in an account at the end of 10 years if $18,000 is deposited at 4% interest compounded semi-annually? Round to the nearest dollar. (Assume no withdrawals are made.)

OpenStudy (callisto):

Amount = P x (1+r%)^n \(= 18000 \times (1+ \frac{0.04}{2})^{10\times 2}\)

OpenStudy (anonymous):

thanks!

OpenStudy (callisto):

Welcome